OpenStudy (anonymous):

Use basic identities to simplify the expression. sin^2 θ + tan^2 θ + cos^ 2 θ sec^2 θ cos^3 θ sin θ tan^2 θ

OpenStudy (anonymous):

those are the answers up there and the first one is the equation

OpenStudy (anonymous):

sin^2 θ+cos^2 θ=1 1+tan^2 θ=1+(sin^2 θ/cos^2 θ)= 1/cos^2 θ=sec^2 θ do you get it?

OpenStudy (anonymous):

why did you put one over tan? Otherwise- i do understand thanks so much :)

OpenStudy (anonymous):

my english isn't enough adequate to explain, dude. excuse me :)